OpenStudy (anonymous):

solve for t. r=2/3t+v

OpenStudy (anonymous):

Subtract v from both sides to get r-v=2/3t. Then multiply by 3/2 to solve for t , giving you (3/2)(r-v)=t.

OpenStudy (anonymous):

Yes it is in my Algebra textbook

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You are welcome (: